Inc. a U.S. based exporter that exclusively sells its products in Great Britain.,Inc expects to sell 2,000 units at a price of GBP 10.00 per unit. inc currently produces in KY and has a variable production cost of USD 8.10 per unit, and total fixed costs of USD 5,000. Use an expected XUSD/GBP = 1.50. What is inc’s expected cash flow from operations?
Answer #1

rate positively ... let me know if you need any clarification..

Computation of expected cash flow from operation
A sales =2000*10*1.5 30000
Less
Cost =2000*8.1 16200
Fixed cost = 5000
B Total cost 21200
C=A-B Cash from operation = 8800